Question 16(5 points) You have just measured a metal cylinder; listed below is a partial table of your data. Fill in the blanks Mass-5.8709 g Diameter 0.53 cm Height-4.40 cm Using the formulas below determine: (V-r2 h, where r radius, h- height, n 3.14) Volume of the cylinder (V)
Answer #1

volume = \pi * r^2 * h

= r = d/2 = 0.53 / 2 = 0.265 cm

h = 4.40 cm

volume = 3.14*(0.265)^2*4.40

=0.9702286 cm^3 or cc

any query please comment

if satisfied please rate it thanks